Last year ABC Corp. had sales of $525,000 and a net income of $12,600, and its year-end assets were $100,000. The firm's total-liabilities-to-total-assets ratio was 50.00%. Based on the DuPont equation, what was ABC's ROE? Show your answer in this format: 12.34%
Answer #1

total-liabilities-to-total-assets ratio=Total liabilities/Total asset

Total liabilities=(100,000*50%)=$50,000

Total assets=Total liabilities+Total equity

Total equity=100,000-50,000=$50,000

Profit margin=Net income/Sales

=(12600/525,000)=0.024

Total asset turnover=Sales/Total assets

=(525,000/100,000)=5.25

Equity multiplier=Total assets/Total equity

=100,000/50,000=2

ROE=Profit margin*Total asset turnover*Equity multiplier

=0.024*5.25*2

=25.2%